OpenStudy (anonymous):

Solve 2x + 5y = –13 3x – 4y = –8 A. (4, 1) B. (–4, 1) C. (4, –1) D. (–4, –1)

OpenStudy (anonymous):

-3(2x+5y=-13) 2(3x-4y=-8) -6x-15y=39 + 6x-8y=-16 --------------------- -23y=23 y=-1 3x+4=-8 3x=-12 x=-4